### v2

@@ 2.3 Sample Element Types
The most basic element considered in the comparison is the original 2-node ANCF beam element 3243, which is fully parameterized, as referenced in gerstmayrECCOMAS2005 . As illustrated in Fig. 2.1, each node is described by 12 coordinates, including a position vector $\mathbf{r}_{i}$ and its gradients with respect to the local coordinates $u$, $v$, and $w$, denoted as $\frac{\partial\mathbf{r}_{i}}{\partial u}$, $\frac{\partial\mathbf{r}_{i}}{\partial v}$, and $\frac{\partial\mathbf{r}_{i}}{\partial w}$. This full set of coordinates per node defines the “fully parameterized” terminology. With two such nodes, the element possesses 24 degrees of freedom when unconstrained. The visualization of this element can be seen in Figure 2.
@@
$=[$ $\mathbf{r}_{1}\;\;\mathbf{r}_{1,u}\;\;\mathbf{r}_{1,v}\;\;\mathbf{r}_{1,w}\;\;$ (2a)
$\mathbf{r}_{2}\;\;\mathbf{r}_{2,u}\;\;\mathbf{r}_{2,v}\;\;\mathbf{r}_{2,w}\,]$
$\equiv[$ $\mathbf{e}_{1}\;\;\mathbf{e}_{2}\;\;\mathbf{e}_{3}\;\;\mathbf{e}_{4}\;\;$
$\mathbf{e}_{5}\;\;\mathbf{e}_{6}\;\;\mathbf{e}_{7}\;\;\mathbf{e}_{8}\,]\;,$
@@
2.3.2 ANCF 3443 Shell Element
@@
2.3.3 10-Node Tetrahedron Element
Figure 4: Visualization of a 10-node tetrahedron (T10) element.
The 10-node tetrahedron (T10), as shown in Figure 4, is a quadratic isoparametric solid element defined on a reference (parent) tetrahedron in the parametric coordinates $(\xi,\eta,\zeta)$ with
$\xi\geq 0,\quad\eta\geq 0,\quad\zeta\geq 0,\quad\xi+\eta+\zeta\leq 1.$ (4)
It contains four corner nodes at the vertices of the simplex and six nodes located at the midpoints of the edges. The nodal positions in the parent domain can be expressed in barycentric coordinates $L_{i}$ ($i=1,\dots,4$),
$L_{1}=1-\xi-\eta-\zeta,\qquad L_{2}=\xi,\qquad L_{3}=\eta,\qquad L_{4}=\zeta,$ (5)
which satisfy $L_{i}\geq 0$ and $L_{1}+L_{2}+L_{3}+L_{4}=1$. The four vertex nodes correspond to $L_{i}=1$ at each corner, and the six mid-edge nodes correspond to $L_{i}=L_{j}=1/2$ on each edge $(i,j)$.
The T10 element employs the quadratic Lagrange basis on the tetrahedral simplex. In terms of barycentric coordinates, the shape functions for the four corner nodes are
$N_{1}$ $=L_{1}(2L_{1}-1),$ (6)
$N_{2}$ $=L_{2}(2L_{2}-1),$
$N_{3}$ $=L_{3}(2L_{3}-1),$
$N_{4}$ $=L_{4}(2L_{4}-1),$
and the six edge-midpoint nodes are
$N_{5}$ $=4L_{1}L_{2},\qquad N_{6}$ $=4L_{2}L_{3},\qquad N_{7}$ $=4L_{3}L_{1},$ (7)
$N_{8}$ $=4L_{1}L_{4},\qquad N_{9}$ $=4L_{2}L_{4},\qquad N_{10}$ $=4L_{3}L_{4}.$
These functions form a complete basis for all polynomials of total degree $\leq 2$ on the tetrahedron (dimension 10) and satisfy the usual isoparametric properties:
$\sum_{a=1}^{10}N_{a}(\xi,\eta,\zeta)=1,\qquad N_{a}(\xi_{b},\eta_{b},\zeta_{b})=\delta_{ab},$ (8)
which implies $C^{0}$ continuity across conforming meshes and Kronecker delta interpolation of nodal degrees of freedom.
Let $\mathbf{X}_{a}$ and $\mathbf{r}_{a}$ denote the nodal positions in the reference and current configurations, respectively, and let $\mathbf{d}_{a}=\mathbf{r}_{a}-\mathbf{X}_{a}$ be the displacement of node $a$. The isoparametric mappings for the reference position and displacement fields inside the element are
$\mathbf{X}(\xi,\eta,\zeta)$ $=\sum_{a=1}^{10}N_{a}(\xi,\eta,\zeta)\,\mathbf{X}_{a},$ (9)
$\mathbf{u}(\xi,\eta,\zeta)$ $=\sum_{a=1}^{10}N_{a}(\xi,\eta,\zeta)\,\mathbf{d}_{a}.$
The current position is $\mathbf{r}=\mathbf{X}+\mathbf{u}$, and the deformation gradient in the Total Lagrangian setting is obtained from
$\mathbf{F}(\xi,\eta,\zeta)=\frac{\partial\mathbf{r}}{\partial\mathbf{X}}=\frac{\partial\mathbf{r}}{\partial(\xi,\eta,\zeta)}\left(\frac{\partial\mathbf{X}}{\partial(\xi,\eta,\zeta)}\right)^{-1}.$ (10)
Both Jacobians are computed from the shape function derivatives $N_{a,\xi}$, $N_{a,\eta}$, $N_{a,\zeta}$, and the nodal coordinates $\mathbf{X}_{a}$, so that the standard Total Lagrangian strain measures and element strain–displacement matrix $\mathbf{B}$ follow directly.
Element integrals in the reference configuration are evaluated over the parent tetrahedron using Gaussian quadrature. In this work we employ a five-point rule that is exact for polynomials of total degree up to three, with quadrature points $(\xi_{q},\eta_{q},\zeta_{q})$ and weights $W_{q}$ chosen such that
$\int_{\Omega_{0}^{e}}(\cdot)\,\mathrm{d}V=\int_{\hat{\Omega}}(\cdot)\,J(\xi,\eta,\zeta)\,\mathrm{d}\xi\,\mathrm{d}\eta\,\mathrm{d}\zeta\approx\sum_{q=1}^{5}(\cdot)_{q}\,J_{q}\,W_{q},$ (11)
where $\hat{\Omega}$ is the parent tetrahedron, $J=\det\big(\partial\mathbf{X}/\partial(\xi,\eta,\zeta)\big)$ is the geometric Jacobian, and $(\cdot)_{q}$ denotes the integrand evaluated at quadrature point $q$. The same quadrature rule is used for both mass and internal-force contributions, providing a consistent and robust discretization for the T10 element.
3 Augmented Lagrangian Formulation
Start by assuming the position-level kinematic constraint equations read:
$c_{i}(\mathbf{r}_{n+1},\mathbf{A}_{n+1},t_{n+1})=0\;,\quad i=1,2,\ldots,m\;,$ (12)
and they combine to form the vector of constraints $\mathbf{c}(\mathbf{r}_{n+1},\mathbf{A}_{n+1},t_{n+1})=\mathbf{0}$. For compactness, define $c_{i}^{n+1}\coloneqq c_{i}(\mathbf{r}_{n+1},\mathbf{A}_{n+1},t_{n+1})$. ALM starts with the Lagrangian function:
$\mathcal{L}_{\bm{\rho}}(\dot{\mathbf{v}}_{n+1},\bar{\bm{\omega}}_{n+1},\bm{\lambda})=\ell(\dot{\mathbf{v}}_{n+1},\bar{\bm{\omega}}_{n+1})+\sum_{i=1}^{m}\Bigl[\lambda_{i}c_{i}^{n+1}+\frac{\rho_{i}}{2}\left(c_{i}^{n+1}\right)^{2}\Bigr]\;,$ (13)
where $\rho_{i}$ are penalty parameters that control the weight of the quadratic penalty term, and $\lambda_{i}$ are the Lagrange multipliers, i.e., the dual variables. The constraint functions $c_{i}(\mathbf{r}_{n+1},\mathbf{A}_{n+1},t_{n+1})$ are evaluated at the position level and depend implicitly on the optimization variables $\dot{\mathbf{v}}_{n+1}$ and $\bar{\bm{\omega}}_{n+1}$ through the integration scheme used to compute the updated position $\mathbf{r}_{n+1}$ and orientation $\mathbf{A}_{n+1}$.
The optimization approach, either first-order or second-order, can be used to minimize the cost functions in Eqs. 13 and 14. For the velocity level approach to solving the constrained multibody dynamics problem, see Eq. 14, the constraints are abstracted as
$\mathbf{c}(\mathbf{x})=\mathbf{0}\;,\quad\text{where }\mathbf{x}=\begin{bmatrix}\dot{\mathbf{v}}_{n+1}\\ \bar{\bm{\omega}}_{n+1}\end{bmatrix}\;$.
The Augmented Lagrangian for this problem is
$\mathcal{L}_{\bm{\rho}}(\mathbf{x},\bm{\lambda})=\ell(\mathbf{x})+\sum_{i=1}^{m}\left[\lambda_{i}\tilde{\nu}_{i}(\mathbf{x})+\frac{\rho_{i}}{2}\tilde{\nu}_{i}(\mathbf{x})^{2}\right]\;,$ (14)
where $\ell(\mathbf{x})$ is the cost function, $\lambda_{i}$ are the Lagrange multipliers, and $\rho_{i}>0$ are penalty weights associated with each constraint.
In the proposed framework for solving the Total Lagrangian finite element analysis for multibody dynamics, we derive the equations of motion from the principle of virtual work. The total virtual work is written as:
$\delta W=\delta W_{\mathrm{inertia}}+\delta W_{\mathrm{applied}}+\delta W_{\mathrm{force\text{-}field}}+\delta W_{\mathrm{internal}}\;,$ (15)
where $\delta W_{\mathrm{inertia}}$, $\delta W_{\mathrm{applied}}$, $\delta W_{\mathrm{force\text{-}field}}$, and $\delta W_{\mathrm{internal}}$ denote the virtual works of the inertial terms, applied forces, mass-distributed force fields, and internal forces, respectively. Imposing $\delta W=0$ and factoring out the virtual generalized displacements yields an expression of the form:
$\sum_{i=1}^{n_{u}}\delta\mathbf{q}_{i}^{T}\,\mathbf{g}_{i}=0\;,$ (16)
where $\delta\mathbf{q}_{i}$ is an arbitrary virtual displacement associated with the $i$th nodal unknown and $\mathbf{g}_{i}$ is the corresponding residual contribution. Since $\delta\mathbf{q}_{i}$ are arbitrary, Eq. (16) implies $\mathbf{g}_{i}=\mathbf{0}$ for all $i$, which yields the discrete equations of motion.
In the following sections, we derive the expressions for all contribution terms that make up Eq. (15). The inertial term is derived in Section 4. The mass-distributed force fields are discussed in Section 5, and the internal forces are discussed in Section 6. Finally, the applied forces, which in the majority of problems come from friction and contact and external applied force loads, are discussed in Section 7.
With the update $\mathbf{q}_{n+1}=\mathbf{q}_{n}+h\,\mathbf{v}_{n+1}$, the residual can be written as:
$\mathbf{g}(\mathbf{v},\bm{\lambda})$ $=\frac{1}{h}\mathbf{M}(\mathbf{v}-\mathbf{v}_{n})+\mathbf{f}_{\mathrm{int}}(\mathbf{q}_{n}+h\mathbf{v},\mathbf{v})-\mathbf{f}_{\mathrm{ext}}-\mathbf{f}_{\mathrm{ff}}$ (17)
$\quad+h\,\mathbf{C}_{q}(\mathbf{q}_{n}+h\mathbf{v})^{T}\bigl(\bm{\lambda}+\rho\,\mathbf{c}(\mathbf{q}_{n}+h\mathbf{v})\bigr)\;,$
where $\mathbf{M}$ is a preassembled and constant mass matrix of the dynamic system, $\rho$ is the ALM penalty parameter, and $\mathbf{c}(\mathbf{q})=0$ denotes bilateral constraints with constraint Jacobian $\mathbf{C}_{q}=\partial\mathbf{c}/\partial\mathbf{q}$. Note that $\mathbf{c}(\mathbf{q})\in\mathbb{R}^{m}$ is a constraint-space quantity and is mapped to the generalized space through $\mathbf{C}_{q}(\cdot)^{T}$ in Eq. (17). We evaluate $\mathbf{C}_{q}^{T}(\cdot)$ via a sparse representation of $\mathbf{C}_{q}^{T}$, and enforce constraints with an augmented-Lagrangian outer loop that alternates an inner solve for $\mathbf{v}$ with multiplier updates until $\|\mathbf{c}\|_{2}$ meets a prescribed tolerance.
In Eq. (17), $\mathbf{f}_{\mathrm{int}}(\mathbf{q},\mathbf{v})$ denotes the assembled internal force vector induced by material deformation (including elastic and viscous contributions), while $\mathbf{f}_{\mathrm{ext}}$ collects applied loads specified directly as external nodal forces (including point loads and contact/friction forces treated as external). The term $\mathbf{f}_{\mathrm{ff}}$ accounts for mass-distributed force fields (body forces) specified per unit mass, such as gravity; although it can be absorbed into $\mathbf{f}_{\mathrm{ext}}$, we keep it separate to emphasize its origin in the force-field virtual work component in Eq. (15).
The residual form in Eq. (17) can be interpreted as the first-order optimality condition of a scalar augmented objective in the velocity unknown. Introduce the step map $\mathbf{q}(\mathbf{v})\coloneqq\mathbf{q}_{n}+h\,\mathbf{v}$, and assume that the internal force operator admits an incremental potential $\Pi_{\mathrm{int}}(\mathbf{q},\mathbf{v})$ such that $\nabla_{\mathbf{v}}\!\bigl(\frac{1}{h}\Pi_{\mathrm{int}}(\mathbf{q}(\mathbf{v}),\mathbf{v})\bigr)=\mathbf{f}_{\mathrm{int}}(\mathbf{q}(\mathbf{v}),\mathbf{v})$. Then the residual $\mathbf{g}(\mathbf{v},\bm{\lambda})$ is obtained by differentiating the discrete augmented Lagrangian cost:
$\Phi_{\rho}(\mathbf{v},\bm{\lambda})$ $=\frac{1}{2h}\,(\mathbf{v}-\mathbf{v}_{n})^{T}\mathbf{M}(\mathbf{v}-\mathbf{v}_{n})$ (18)
$\quad+\frac{1}{h}\,\Pi_{\mathrm{int}}(\mathbf{q}_{n}+h\mathbf{v},\mathbf{v})-\mathbf{f}_{\mathrm{ext}}^{T}\mathbf{v}-\mathbf{f}_{\mathrm{ff}}^{T}\mathbf{v}$
$\quad+\bm{\lambda}^{T}\mathbf{c}(\mathbf{q}_{n}+h\mathbf{v})$
$\quad+\frac{\rho}{2}\,\bigl\|\mathbf{c}(\mathbf{q}_{n}+h\mathbf{v})\bigr\|_{2}^{2}.$
Using $\nabla_{\mathbf{v}}\mathbf{c}(\mathbf{q}_{n}+h\mathbf{v})=h\,\mathbf{C}_{q}(\mathbf{q}_{n}+h\mathbf{v})$ and the chain rule, one verifies that $\nabla_{\mathbf{v}}\Phi_{\rho}(\mathbf{v},\bm{\lambda})=\mathbf{g}(\mathbf{v},\bm{\lambda})$. Consequently, for fixed multipliers $\bm{\lambda}$, the inner solve $\mathbf{g}(\mathbf{v},\bm{\lambda})=\mathbf{0}$ is equivalent to enforcing stationarity of the augmented cost (18) with respect to $\mathbf{v}$, while the outer ALM loop updates $\bm{\lambda}$ to reduce the constraint violation $\|\mathbf{c}\|_{2}$.
4 Inertia Force Contribution in TL-FEA Formulation
The inertia contribution in Eq. (15) is
where $\rho_{r}(\mathbf{u})$ is the mass density expressed in the reference configuration and $\mathbf{u}=[u,v,w]^{T}$ are the material coordinates. Using the TL interpolation $\mathbf{r}(\mathbf{u};t)=\sum_{i=1}^{n_{u}}\mathbf{e}_{i}(t)s_{i}(\mathbf{u})$, we obtain
$\delta W_{\mathrm{inertia}}=-\sum_{i=1}^{n_{u}}\delta\mathbf{e}_{i}^{T}\sum_{j=1}^{n_{u}}m_{ij}\,\ddot{\mathbf{e}}_{j}\;,\qquad m_{ij}=\int_{V_{r}}\rho_{r}(\mathbf{u})\,s_{i}(\mathbf{u})\,s_{j}(\mathbf{u})\,\mathrm{d}V_{r}\;.$ (20)
The coefficients $m_{ij}=m_{ji}$ depend only on reference quantities ($\rho_{r}$ and the shape functions), hence they are constant and can be precomputed.
It is convenient to introduce the element unknown vector
$\mathbf{e}=[\mathbf{e}_{1}^{T},\ldots,\mathbf{e}_{n_{u}}^{T}]^{T}\in\mathbb{R}^{3n_{u}}\;.$
Equation (20) can then be written as
$\delta W_{\mathrm{inertia}}=-\delta\mathbf{e}^{T}\mathbf{M}_{e}\,\ddot{\mathbf{e}},\qquad\mathbf{M}_{e}=\bigl[m_{ij}\mathbf{I}_{3}\bigr]_{i,j=1}^{n_{u}}\;.$ (21)
After standard FE assembly over all elements, Eq. (21) yields the global inertial virtual work $\delta W_{\mathrm{inertia}}=-\delta\mathbf{q}^{T}\mathbf{M}\,\ddot{\mathbf{q}}$, where $\mathbf{M}$ is the global consistent mass matrix. In particular, under the backward-Euler update used to obtain Eq. (17), $\ddot{\mathbf{q}}_{n+1}=\dot{\mathbf{v}}_{n+1}\approx(\mathbf{v}_{n+1}-\mathbf{v}_{n})/h$, which leads to the inertial residual contribution $\frac{1}{h}\mathbf{M}(\mathbf{v}-\mathbf{v}_{n})$ in Eq. (17).
5 Mass-Distributed Force Field Contribution in TL-FEA Formulation
The force-field component $\delta W_{\mathrm{force\text{-}field}}$ in Eq. (15) accounts for mass-distributed loads specified as a force per unit mass. Let $\mathbf{b}(P,t)\in\mathbb{R}^{3}$ denote such a field (e.g., gravity) evaluated at the current spatial position $P=\mathbf{r}(\mathbf{u};t)$ of a material point with reference coordinates $\mathbf{u}=[u,v,w]^{T}$. Its virtual work over a reference volume $V_{r}$ is
$\delta W_{\mathrm{force\text{-}field}}=\int_{V_{r}}\rho_{r}(\mathbf{u})\,\delta\mathbf{r}(\mathbf{u};t)^{T}\mathbf{b}(\mathbf{r}(\mathbf{u};t),t)\,\mathrm{d}V_{r}\;,$ (22)
where $\rho_{r}(\mathbf{u})$ is the reference density. Using $\delta\mathbf{r}(\mathbf{u};t)=\sum_{i=1}^{n_{u}}\delta\mathbf{e}_{i}^{T}s_{i}(\mathbf{u})$, Eq. (22) becomes
$\delta W_{\mathrm{force\text{-}field}}=\sum_{i=1}^{n_{u}}\delta\mathbf{e}_{i}^{T}\mathbf{f}^{\mathrm{ff}}_{i},\qquad\mathbf{f}^{\mathrm{ff}}_{i}\coloneqq\int_{V_{r}}\rho_{r}(\mathbf{u})\,s_{i}(\mathbf{u})\,\mathbf{b}(\mathbf{r}(\mathbf{u};t),t)\,\mathrm{d}V_{r}\in\mathbb{R}^{3}.$ (23)
After element assembly, the force-field contribution is collected in $\mathbf{f}_{\mathrm{ff}}$ in Eq. (17).
$\mathbf{f}^{\mathrm{ff}}_{i}=\rho_{r}\,\mathbf{b}_{0}\,v_{i},\qquad v_{i}\coloneqq\int_{V_{r}}s_{i}(\mathbf{u})\,\mathrm{d}V_{r}\;.$ (24)
Note that, although the integration is carried out over $V_{r}$, the field must be evaluated at the current position $P=\mathbf{r}(\mathbf{u};t)$; consequently, $\mathbf{b}$ can vary in time as the body moves through space.
6 Internal Force Contribution in TL-FEA Formulation
The internal force contribution to the equations of motion (EOM) in the Total Lagrangian finite element analysis (TL-FEA) formulation arises from the material’s constitutive response to deformation. In the proposed framework, the internal forces are the result of two major components: the hyperelastic material response and the viscous damping effect. The former captures the elastic behavior of the material under finite deformation, while the latter provides numerical stability and models energy dissipation during dynamic simulations.
6.1 Hyperelastic Material Laws
6.1.1 Generic TL internal force expression
The contributions of the internal forces to the EOM for the ANCF beam element and the ANCF 3443 shell element are discussed and derived as follows:
Following the notation used in Eq. 2a, the deformation gradient for the ANCF 3243 element is:
$\mathbf{F}$ $=\frac{\partial\mathbf{r}}{\partial\mathbf{u}}=\mathbf{N}(t)\cdot\frac{\partial\mathbf{s}}{\partial\mathbf{u}}$
$=\mathbf{N}(t)\cdot\left[\begin{smallmatrix}s_{1,\mathbf{u}}\\ s_{2,\mathbf{u}}\\ s_{3,\mathbf{u}}\\ s_{4,\mathbf{u}}\\ s_{5,\mathbf{u}}\\ s_{6,\mathbf{u}}\\ s_{7,\mathbf{u}}\\ s_{8,\mathbf{u}}\end{smallmatrix}\right]=\mathbf{e}_{1}s_{1,\mathbf{u}}+\mathbf{e}_{2}s_{2,\mathbf{u}}+\ldots+\mathbf{e}_{8}s_{8,\mathbf{u}}$
$=\sum_{i=1}^{8}\mathbf{e}_{i}s_{i,\mathbf{u}}\in\mathbb{R}^{3\times 3}\;.$
Following the notation used earlier, the deformation gradient $\mathbf{F}$ for the 3443 shell element is:
$\mathbf{F}=\frac{\partial\mathbf{r}}{\partial\mathbf{u}}=\sum_{i=1}^{16}\mathbf{e}_{i}\,s_{i,\mathbf{u}}\in\mathbb{R}^{3\times 3}$
where each $s_{i,\mathbf{u}}\in\mathbb{R}^{1\times 3}$.
6.1.2 St Venant–Kirchhoff (SVK) material
The Saint-Venant–Kirchhoff (SVK) material model is a fundamental hyperelastic constitutive model in finite element analysis (FEA), particularly suited to the total Lagrangian formulation, where it employs the Green–Lagrange strain tensor and the second Piola–Kirchhoff stress to capture moderate nonlinear deformations in compressible solids Bonet2008nonlinear . This model extends classical linear elasticity to finite strains by defining a quadratic strain energy density function $\Psi(\mathbf{E})$ (see (25)), with $\lambda$ and $\mu$ denoting the Lamé constants, making it computationally straightforward for quasi-static and dynamic simulations of structures undergoing large rotations but limited stretches Bonet2022limitations . Popular in computational mechanics for its simplicity and compatibility with isoparametric elements, the SVK model finds broad use in beam and shell analyses, such as Timoshenko–Ehrenfest beams, biological soft tissues, and multifunctional laminates, although its popularity wanes in extreme compression due to unphysical softening beyond about $58\%$ strain Zur2023large . Despite its limitations in large uniaxial compression, it remains a standard benchmark for validating more advanced hyperelastic models, such as neo-Hookean or Mooney–Rivlin formulations, in FEA software like Abaqus Luo2020benchmark .

### v3

@@ A.3 The 3243 ANCF Beam Element
The most basic element considered in the comparison is the original 2-node ANCF beam element 3243, which is fully parameterized, as referenced in [37]. As illustrated in Fig. 4, each node is described by 12 coordinates, including a position vector $\mathbf{r}_{i}$ and its gradients with respect to the local coordinates $u$, $v$, and $w$, denoted as $\frac{\partial\mathbf{r}_{i}}{\partial u}$, $\frac{\partial\mathbf{r}_{i}}{\partial v}$, and $\frac{\partial\mathbf{r}_{i}}{\partial w}$. This full set of coordinates per node defines the “fully parameterized” terminology. With two such nodes, the element possesses 24 degrees of freedom when unconstrained. The visualization of this element can be seen in Figure 4.
@@
$=[$ $\mathbf{r}_{1}\;\;\mathbf{r}_{1,u}\;\;\mathbf{r}_{1,v}\;\;\mathbf{r}_{1,w}\;\;\mathbf{r}_{2}\;\;\mathbf{r}_{2,u}\;\;\mathbf{r}_{2,v}\;\;\mathbf{r}_{2,w}\,]$ (91a)
$\equiv[$ $\mathbf{e}_{1}\;\;\mathbf{e}_{2}\;\;\mathbf{e}_{3}\;\;\mathbf{e}_{4}\;\;\mathbf{e}_{5}\;\;\mathbf{e}_{6}\;\;\mathbf{e}_{7}\;\;\mathbf{e}_{8}\,]\;,$
@@
A.4 ANCF 3443 Shell Element
@@
Appendix B Joint Constraint Details
This appendix records the lower-level details behind the dot-product constraints introduced in Subsec.˜2.2.5. The accumulated revolute-joint Jacobian blocks, their insertion into the augmented-Lagrangian residual, and the main consequences for the Newton system are now discussed in the main text. What remains here are the detailed scaling argument behind row normalization and the explicit block form of the DP1 Hessian.
B.1 Scaling Basis for Constraint Row Normalization
A mixed CD/DP1 constraint set introduces a magnitude mismatch that can severely degrade Newton convergence. CD constraint values are $\mathcal{O}(\|\Delta\mathbf{r}\|)$, whereas DP1 values scale as $\mathcal{O}(\delta^{2})$ when the direction vectors are defined by material fibers of length $\delta$, chosen as a small fraction of the element diameter. The Jacobian entries inherit the same scaling: $\mathcal{O}(1)$ for CD and $\mathcal{O}(\delta)$ for DP1.
Two consequences follow. First, the raw Newton residual barely corrects angular errors because the DP1 rows are much smaller than the CD rows. Second, in the Gauss–Newton term $h^{2}\rho\,\mathbf{C}_{q}^{\mathsf{T}}\mathbf{C}_{q}$, the angular eigenvalues are $\mathcal{O}(\rho\delta^{2})$, whereas the translational eigenvalues are $\mathcal{O}(\rho)$. The condition number is therefore amplified by a factor of order $1/\delta^{2}$. This is a conditioning artifact introduced by the choice of short material fibers used to define the joint directions; it is not a property of the underlying mechanics.
The remedy adopted in Eqs.˜25 and 26 is per-row normalization with constant weights computed once from the reference configuration. For the revolute joint, the three CD rows are left unscaled, while each DP1 row is divided by the reference magnitude of the two participating directions. When $|\mathbf{a}_{0}|=|\mathbf{b}_{j,0}|=\delta$, the DP1 weights are of order $1/\delta$, so the normalized DP1 rows become $\mathcal{O}(1)$ at the level of both residual and Jacobian. After normalization, the translational and angular parts of the Gauss–Newton term are both $\mathcal{O}(\rho)$, which restores a balanced Newton system.
The same argument extends to any joint containing dot-product constraints. The precise weights may differ from one joint type to another, but the principle is unchanged: the constraint rows should be scaled so that their Jacobian magnitudes are commensurate before entering the Newton solve.
B.2 Explicit DP1 Hessian Blocks
For the revolute joint of Subsec.˜2.2.5, let the point set be $(\mathcal{P},\mathcal{Q})$ on body $b$ with hosting elements $(E,F)$ and $(\mathcal{R},\mathcal{S},\mathcal{T})$ on body $c$ with hosting elements $(G,H,K)$. The evaluation operators are those introduced in Eq.˜22. Consider first the DP1 constraint
$c_{\mathrm{DP1},1}=\mathbf{a}^{\mathsf{T}}\mathbf{b}_{1}-f_{1},$ (93)
with $\mathbf{a}=\mathbf{r}_{\mathcal{Q}}-\mathbf{r}_{\mathcal{P}}$ and $\mathbf{b}_{1}=\mathbf{r}_{\mathcal{S}}-\mathbf{r}_{\mathcal{R}}$. Since $c_{\mathrm{DP1},1}$ is a product of two functions that are linear in $\mathbf{q}$, its Hessian is constant. Ordered as $(\mathbf{e}_{E}^{b},\mathbf{e}_{F}^{b},\mathbf{e}_{G}^{c},\mathbf{e}_{H}^{c})$, the nonzero block structure is
$\nabla_{q}^{2}c_{\mathrm{DP1},1}=\begin{bmatrix}\mathbf{0}&\mathbf{0}&+\bm{\Sigma}_{E}^{\mathsf{T}}\bm{\Sigma}_{G}&-\bm{\Sigma}_{E}^{\mathsf{T}}\bm{\Sigma}_{H}\\[4.0pt] \mathbf{0}&\mathbf{0}&-\bm{\Sigma}_{F}^{\mathsf{T}}\bm{\Sigma}_{G}&+\bm{\Sigma}_{F}^{\mathsf{T}}\bm{\Sigma}_{H}\\[4.0pt] +\bm{\Sigma}_{G}^{\mathsf{T}}\bm{\Sigma}_{E}&-\bm{\Sigma}_{G}^{\mathsf{T}}\bm{\Sigma}_{F}&\mathbf{0}&\mathbf{0}\\[4.0pt] -\bm{\Sigma}_{H}^{\mathsf{T}}\bm{\Sigma}_{E}&+\bm{\Sigma}_{H}^{\mathsf{T}}\bm{\Sigma}_{F}&\mathbf{0}&\mathbf{0}\end{bmatrix}$. (94)
For the second DP1 row, $c_{\mathrm{DP1},2}=\mathbf{a}^{\mathsf{T}}\mathbf{b}_{2}-f_{2}$, the structure is identical, with $\bm{\Sigma}_{H}$ replaced by $\bm{\Sigma}_{K}$.
Three properties are worth recording. First, the Hessian is constant: it depends only on shape-function evaluations at the chosen constraint points and can therefore be precomputed once the joint topology is fixed. Second, it is sparse: only cross-body couplings appear, while same-body diagonal blocks vanish. Third, it is indefinite: the vanishing diagonal blocks together with the nonzero off-diagonal couplings imply eigenvalues of both signs.
These properties explain the structure of the Newton matrix in Eqs.˜28 and 29. The Gauss–Newton contribution is positive semidefinite, whereas the curvature contribution associated with DP1 rows is indefinite and weighted by the effective multipliers $\eta_{i}$. Consequently, once the consistent DP1 Hessian is retained, the full Newton system is symmetric but not, in general, positive definite. This is the algebraic reason that dot-product constraints require a symmetric-indefinite solve if one wishes to preserve the consistent quadratic-convergence model.
Appendix C Select Material Models
C.1 Hyperelastic Material Laws
C.1.1 St Venant-Kirchhoff (SVK) material
The Saint-Venant–Kirchhoff (SVK) material model is a fundamental hyperelastic constitutive model in finite element analysis (FEA), particularly suited to the total Lagrangian formulation, where it employs the Green–Lagrange strain tensor and the second Piola–Kirchhoff stress to capture moderate nonlinear deformations in compressible solids [13]. This model extends classical linear elasticity to finite strains by defining a quadratic strain energy density function $\Psi(\mathbf{E})$ (see (95)), with $\lambda$ and $\mu$ denoting the Lamé constants, making it computationally straightforward for quasi-static and dynamic simulations of structures undergoing large rotations but limited stretches [38]. Popular in computational mechanics for its simplicity and compatibility with isoparametric elements, the SVK model finds broad use in beam and shell analyses, such as Timoshenko–Ehrenfest beams, biological soft tissues, and multifunctional laminates, although its popularity wanes in extreme compression due to unphysical softening beyond about $58\%$ strain [39]. Despite its limitations in large uniaxial compression, it remains a standard benchmark for validating more advanced hyperelastic models, such as neo-Hookean or Mooney–Rivlin formulations, in FEA software like Abaqus [40].